A car's speed changes from 15m/s to 25m/s in 5.0s along a straight road. Find the
acceleration of the car during this period.
1
Expert's answer
2020-09-01T07:10:49-0400

The acceleration of the car, when it changes the speed from "v_0" to "v_1" in time "T" is:"a = \frac{v_1 - v_0}{T} = \frac{25 \frac{m}{s} - 15 \frac{m}{s}}{5 s} = 2 \frac{m}{s^2}".